You have just been hired as a financial analyst for Lydex Company, a manufacturer of safety helmets. Your boss has asked youTo begin your assignment you gather the following financial data and ratios that are typical of companies in Lydex CompanysCompanys industry: Current ratio Acid-test ratio Average collection period Average sale period Return on assets Debt-to-equi
Answer #1

2.

a) Earning per share= Net Income/No. of shares outstanding

This Year last Year
Net Income(A) 630,000 1,167,600
No of Shares(B) 100,000 100,000
Earning per share(C)=(A)/(B) 6.3 11.676

b) Dividend yield Ratio= Dividend per share/Market price per share

This Year last Year
Dividend(A) 530,000 583,800
No of Shares(B) 100,000 100,000
Dividend per share(C)=(A)/(B) 5.3 5.838
Market price per Share(D) 112 80
Dividend yield Ratio=(C)/(D) 4.7% 7.3%

c) Dividend payout ratio= Dividend per share/Earning per share

This Year last Year
Dividend per share(A) 5.3 5.838
Earning per share(B) 6.3 11.676
Dividend payout ratio(C)=(A)/(B) 84.1% 0.5%

d) Price earning ratio= Market price per Share/ Earning Per share

This Year last Year
Market price per Share(A) 112 80
Earning per share(B) 6.3 11.676
Price earning ratio(C)=(A)/(B) 17.78 6.85

e) Book Value per share= Total Shareholder's Equity/ No. of shares outstanding

This Year last Year
Total Shareholders' Equity (A) 9,330,000 9,230,000
No of Shares(B) 100,000 100,000
Book Value per share(C)=(A)/(B) 93.3 92.3

3.

a) Working Capital= Current Assets-Current Liabilities

This Year last Year
Current Assets(A) 7,750,000 6,370,000
Current Liabilities(B) 4,400,000 3,160,000
Working Capital(A)-(B) 3,350,000 3,210,000

b) The current Ratio= Current Liabilities/Current Assets

This Year last Year
Current Assets(A) 7,750,000 6,370,000
Current Liabilities(B) 4,400,000 3,160,000
Current Ratio(A)/(B) 1.76 2.02

c)Acid test Ratio=(Current Assets- Inventory-prepaid Exp)/ Current Liabilities

This Year last Year
Current Assets(A) 7,750,000 6,370,000
Inventory(B) 3,400,000 2,400,000
Prepaid Expenses(C) 240,000 220,000
Quick Assets(D)=(A)-(B)-(C) 4,110,000 3,750,000
Current Liabilities(E) 4,400,000 3,160,000
Acid Test Ratio(A)/(E) 0.93 1.19

d) Average collection period=365/Account Receivable Turnover

This Year last Year
Net credit Sales(A) 15,950,000 14,480,000
Average Account Receivable(B) 2610000 1960000
Account Receivable Turnover Ratio(C)=(A)/(B) 6.11 7.39
Average Collection period=365/(C) 59.74 days 49.39 Days
Account Receivable, Beginning(D) 2160000 1760000
Account Receivable, Ending(E) 3060000 2160000
Average Account Receivable(D)+(E)/2 2,610,000 1960000

e) Average Sale period=365/Inventory Turnover

This Year last Year
Cost of Goods Sold(A) 12760000 10860000
Average Inventories(B) 2900000 2260000
Inventory Turnover Ratio(C)=(A)/(B) 4.4 4.81
Average Sale period=365/(C) 82.95 days 75.88 days
Inventory, Beginning(D) 2,400,000 2120000
Inventory, Ending(E) 3,400,000 2,400,000
Average Inventory(D)+(E)/2 2900000 2260000

f) Operatig Cycle=Average collection period+Average Sale period

This Year last Year
Average Collection period(A) 59.74 days 49.39 Days
Average Sale period(B) 82.95 days 75.88 days
Operatig Cycle(C)=(A)+(B) 142.69 days 125.27 days

g) Total asset Turnover=Net Sales/Average Total Asset

This Year last Year
Net credit Sales(A) 15,950,000 14,480,000
Average Total Assets(B) 16,480,000 14,335,000
total assets Turnover (C)=(A)/(B) 0.97 1.01
Total Assets, Beginning(D) 15,510,000 13160000
Total Assets, Ending(E) 17,450,000 15,510,000
Average Total Assets(D)+(E)/2 16,480,000 14,335,000

please Rate if it helps you or let me know if anything else.